OpenStudy (anonymous):

what percent of 30 is 15

mathslover (mathslover):

let the percent be x : given : x % of 30 = 15

OpenStudy (amistre64):

well, since both are divisible by 5; we can get the same results if we divide and get what % of 6 is 3?

OpenStudy (anonymous):

50%

OpenStudy (amistre64):

good

mathslover (mathslover):

Always remember that : we can write x% as : \(\large{\frac{x}{100}}\) hence \[\large{\frac{x}{100}*30=15}\] \[\large{\frac{x}{\cancel{100}^{50}}*\cancel{30}^{15}=15}\] \[\large{\frac{x}{\cancel{50}^{10}}*\cancel{15}^3=15}\] \[\large{\frac{x}{10}*3=15}\] can u solve for x now ?

mathslover (mathslover):

well this is the same as @amistre64 got u know about that .. do use amistre's method for ease thanks!!!

mathslover (mathslover):

Also : \[\Huge\mathbb{Welcome}\textbf{To}\mathbb{Open}\textbf{Study}\]

mathslover (mathslover):

For more info regarding this site: have a look here : http://openstudy.com/code-of-conduct